# Escalating LLM-based Code Translation Benchmarking into the Class-level Era

### v2

@@ 4. Experimental Design
Category Model Base Model Model Type Time Size #Tokens
Hybrid LLM DeepSeek-V2.5 (DeepSeek Team, 2024a)
DeepSeek-V2-Coder-Instruct (Zhu et al., 2024)
and DeepSeek-V2-Chat (Liu et al., 2024a)
Alignment Model (DeepSeek Team, 2024b) 2024-09 236B 8.1 trillion
General LLM Llama3 (Llama Team, 2024) / Instruction Model 2024-04 8B 15.6 trillion
Llama3 (Llama Team, 2024) / Instruction Model 2024-04 70B 15.6 trillion
Gemma (Team et al., 2024) Gemini (Team et al., 2023) Base Model 2024-02 7B 6 trillion
Code LLM CodeLlama (Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023) Llama2 (Touvron et al., 2023) Instruction Model 2023-08 7B 500 billion
CodeGemma (Team, 2024) Gemma (Team et al., 2024) Base Model 2024-04 7B 500 billion
@@
    In the follow-up sections, we use “DeepSeek” to denote DeepSeek-V2.5 for abbreviation, while using “Llama3-70B” and “Llama3-8B” to discriminate the two versions of Llama3.
@@ 4.1. Studied LLMs
This paper focuses on very recent models released since late 2023. Table 2 lists our studied LLMs with their categories, base models, model type, released times (Time), parameter sizes (Size), and the pretraining token numbers (#Tokens). As can be seen, to explore performance discrepancies among different types, we select LLMs from general (i.e., Llama 3 (Llama Team, 2024) and Gemma (Team et al., 2024)), code (i.e., CodeLlama (Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023) and CodeGemma (Team, 2024)), and hybrid (i.e., DeepSeek-V2.5111We partition DeepSeek-V2.5 to the hybrid category as it is an upgraded version that combines DeepSeek-V2-Chat and DeepSeek-Coder-V2-Instruct. (DeepSeek Team, 2024a)) categories. Besides, for general and code LLMs, we select two for each category to further investigate the differences brought about by training strategies or key techniques. Furthermore, for LLMs from the same family, we also select two different sizes (i.e., Llama3-8B and Llama3-70B) to study the influence of the volume of model parameters.
@@ 4.4. Implementation Details
We utilize six LLMs for our experiments. Specifically, we employ the Deepseek model via its official API interface(Deepseek, 2024). Additionally, CodeGemma, Gemma, and Llama-70B are accessed through the open API interface provided by NVIDIA(NVIDIA, 2024), while Llama-8B and CodeLlama are operated via the open API interface provided by Baidu Qianfan platform(Baidu Qianfan, 2024). The hyperparameters for the code translation process are set as follows: $temperature=0.8$, $top\_p=0.95$, $top\_k=50$, and $n=1$, which means we only fetch LLMs’ first translated candidates for evaluation during inference. All other hyperparameters are kept by default.
Figure 1. Translation results in terms of CSR, CAc, and CAm on benchmarks of Yang et al. and ClassEval-T
@@
DeepSeek 46.81 62.18 94.68 25.53 34.20 64.89 31.91 61.92 96.81 8.51 9.33 77.66 3.19 4.40 79.79 19.15 24.35 51.06
Llama3-70B 36.17 54.15 86.17 4.26 6.48 26.60 24.47 50.52 77.66 5.32 3.89 57.45 5.32 8.29 45.74 7.45 10.10 18.09
Llama3-8B 28.72 51.55 90.43 3.19 5.18 8.51 20.21 48.96 94.68 1.06 2.59 57.45 2.13 3.11 61.70 4.26 3.11 8.51
Gemma 21.28 36.79 84.04 2.13 3.89 8.51 11.70 30.57 70.21 1.06 1.04 22.34 1.06 1.04 25.53 3.19 3.37 7.45
CodeLlama 20.21 39.12 84.04 1.06 1.55 3.19 12.77 35.49 76.60 3.19 3.37 29.79 1.06 2.59 35.11 2.13 2.07 4.26
CodeGemma 26.60 46.63 86.17 1.06 1.30 5.32 20.21 46.37 85.11 5.32 5.18 39.36 3.19 3.63 51.06 5.32 4.92 12.77
Average 29.96 48.40 87.59 6.21 8.77 19.50 20.21 45.64 83.51 4.08 4.23 47.34 2.66 3.84 49.82 6.91 7.99 17.02
@@
    Table cells with light cyan background denote the highest performance in terms of CAc/CAm/CSR on each translation pair among various LLMs. Digits in bold and underline denote the highest and second-highest average performance in terms of CAc/CAm/CSR among various translation pairs.
@@ 5.1. RQ1: Overall Correctness
Apart from the experiments on ClassEval-T, we also evaluate each studied LLM on Yang et al.’s code translation dataset (Yang et al., 2024a), which is cleaned from (Roziere et al., 2020), one of the most widely used method-level datasets in the code translation field (Roziere et al., 2021; Liu et al., 2023; Szafraniec et al., 2023). Figure 1 demonstrates the correctness comparison between code translation benchmarks of ClassEval-T and Yang et al. among diverse LLMs in terms of CA and CSR, where the holistic translation strategy is adopted here, and detailed strategy comparison results are showcased in Section 5.2. As can be seen, both class-level CA (CAc) and method-level CA (CAm) are adopted for assessment on ClassEval-T, while Yang et al.’s dataset is solely evaluated on CAm as it only consists of method-level code translation samples. Table 3 presents specific experimental results with the holistic translation strategy on ClassEval-T. Based on Table 3 and Figure 1, we have the following observations.
Method-level code translation v.s. class-level code translation. As shown in Figure 1, the code translation performance of LLMs degenerates significantly and consistently on all translation pairs when converting from method-level to class-level benchmark. Specifically, all LLMs studied can achieve CAm scores over 70% and CSR scores over 90% on almost all translation pairs in Yang et al.’s benchmark. However, when assessing them on the ClassEval-T benchmark, LLMs’ code translation performance declines 69.50%–82.65% on average in terms of CAm while 21.81%–59.90% on average in terms of CSR, where half of the translation results suffer over 90% decline in terms of CAm, and almost half of them suffer over 50% decline in terms of CSR. Besides, based on Yang et al.’s benchmark, it is hard to discriminate which LLM performs better, as method-level code translation benchmarks only contain standalone functions, they are shorter, lack diverse dependencies and complex contextual information, making LLMs easy to handle such samples and demonstrate inapparent performance discrepancies. Even in domain-specific PLs, such as Rust and Swift, a recent study (Tao et al., 2024) also unveiled a similar result. Therefore, we conclude that standalone functions are nearly no longer effective in assessing LLMs’ code translation performance, but existing LLMs still have limited capability in complex coding tasks that bring about brand-new challenges, such as class-level code translation.
@@
Comparison among LLMs. Focusing on the results of the ClassEval-T benchmark, DeepSeek always performs the best, except when translating from C++ to Java, it performs second, achieving a CAc score of 22.52%, a CAm score of 32.73%, and a CSR score of 77.48% on average among different translation pairs. Because it carries many more parameters than other LLMs (236B v.s. ¿10B) and upgrades from the combination of two Mixture-Of-Experts (MOE)-driven LLMs of general and code versions. Regarding the comparison between general and code LLMs, CodeGemma, a code LLM, always performs better than Gemma, a general LLM, on all evaluation metrics among various translation pairs. However, Llama3-8B performs better than CodeLlama in most of the translation pairs. The reason behind this performance difference is that CodeGemma is enhanced from Gemma with an extra 500 billion code tokens for code-specific pretraining, resulting in its superior performance against Gemma. In contrast, CodeLlama is based on Llama2 (Touvron et al., 2023), while Llama3 benefits from better data quality, more diverse training data, and a larger training scale. These improvements make Llama3 stronger than Llama2 (Dubey et al., 2024), on which CodeLlama is built. The above phenomenon and analysis indicate that code LLMs are not necessarily better than general LLMs, even though they are specially pre-trained on code corpus. Delving deep into each category, CodeGemma constantly outperforms CodeLlama, while Llama3-8B constantly outperforms Gemma in terms of both CA and CSR, which are in accordance with evaluation results in other coding tasks conducted by recent studies (Team, 2024; Dubey et al., 2024). In addition, it is obvious that Llama3-70B always generates more samples that can pass test suites than Llama3-8B does among all translation pairs, even though Llama3-8B can generate more compilable samples in certain translation pairs, illustrating that LLMs with more parameters carry more powerful code translation ability, given other factors the same.
@@
Comparison among PLs. Based on Figure 1 and Table 3, LLMs generally perform much better on Python-oriented translations (e.g., C++/Java-to-Python) in terms of all evaluation metrics. However, concerning the rest of the code translation pairs, LLMs’ performance drops dramatically. Specifically from 20.21%–29.96% to 2.66%–4.08% in terms of CAc, 45.64%–48.40% to 3.84%–4.23% in terms of CAm, and 83.51%–87.59% to 17.02%–19.50% in terms of CSR for C++ and Java-oriented translations. The performance differences can be attributed to pre-training data biases. Recent LLMs and those included in this paper normally use open-sourced code samples on GitHub for coding enhancement pretraining (Zhao et al., 2024; Liu et al., 2024a; Team, 2024; Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023; Tao et al., 2024), while Python has achieved the second most used PL on GitHub since 2022, as reported in (GitHub Octoverse, 2024). Therefore, LLMs potentially gained more knowledge of Python generation, which results in its much superior advantages against C++/Java-oriented translations.
@@ 5.2. RQ2: Translation Strategies
Figure 2 demonstrates the experimental results on ClassEval-T with different code translation strategies mentioned in Section 4.2, where each column showcases experimental results of diverse LLMs in terms of each evaluation metric (either CAc, CAm, or CSR) among six translation pairs. Our observations are listed below.
Analysis among LLMs. As shown in Figure 2, the holistic translation strategy is always the best selection for DeepSeek, outperforming the other two strategies by 64.94%–337.93% in terms of CAc, by 64.07%–172.66% in terms of CAm, and by 65.53%–72.05% in terms of CSR, while for other LLMs, it depends on different translation scenarios. One potential reason might be that DeepSeek carries 236B parameters upgraded from the combination of its previous general and code versions, empowering exceptional code understanding and generation ability. Thus, it is not difficult to complete the code translation task given the entire class as inputs in the holistic strategy. However, other LLMs, with parameter sizes that are smaller than 10B, carry relatively limited capability in leveraging long input context, as revealed by (Liu et al., 2024b). Thus, they may fall short in some scenarios with a holistic code translation strategy but perform better with the min-dependency strategy. Besides, LLMs with the standalone strategy perform the worst on CAc and CAm in most of the translation scenarios, as we do not provide contextual dependencies but a standalone class skeleton or method during the translation, leading to the translated methods or fields being highly possible to have conflicts with other translated counterparts in classes. For example, one method invoked by another may be translated to different method names during the translation of its caller and itself, causing conflicts in a class. However, this is not the case in terms of CSR. When translating with the standalone strategy, LLMs tend to generate some contextual dependencies based on their imagination; thus, discriminating the translated results from other generated contexts via manually crafted rules is one of our processes for the standalone translation strategy. Considering that the translation process of LLMs is uncontrollable, which means they may rename the source program completely, making our rules difficult to capture the translated results. In this paper, we deem the above a failed translation, and no method will be inserted into the translated class skeleton. This will definitely affect functionality correctness but not the program compilation, leading to an overestimation of CSR among LLMs in the standalone translation strategy.
@@
Analysis among PLs. The selection of translation strategy also relies on target PLs. As shown in Figure 2, the holistic strategy always performs the best in Python-oriented translations in terms of CAc and CAm, outperforming other strategies on average by 42.05%–84.40% in terms of CAc while by 35.22%–59.09% in terms of CAm. However, focusing on C++-oriented translation, the min-dependency strategy manifests its superiority, surpassing holistic and standalone strategies by 30.84%–85.05% and 20.65%–72.80% in terms of CAc and CAm, respectively. The reasons are multifaceted: (1) C++ codes in ClassEval-T are longer than Python codes, where the former carries 995.43 tokens on average while the latter contains 705.33 ones, using CodeLlama for tokenization; thus, when translating to C++, LLMs, except DeepSeek (default output limitation=4k tokens), might suffer output limits (all are fixed to 1k tokens by default), leading to some of the classes not being generated completely with the holistic strategy. (2) The min-dependency strategy aligns with C++’s separate compilation feature, enabling flexible code splicing and tolerance for unimplemented sections (Favre, Jean-Marie, 1995). In addition, C++ is more complex than Python and Java, especially in terms of manual memory management, template systems, and strict dependency management. Generating code step by step reduces the information that LLM needs to pay attention to each time, reducing the possibility of LLM generating incorrect code. In this case, the min-dependency strategy becomes more suitable for translating other PLs to C++, while the holistic strategy is recommended for Python-oriented translations. Nonetheless, equivalent programs written in Java also contain a relatively large number of tokens, which is 718.66 tokens per class in ClassEval-T, using CodeLlama as the tokenization. However, it carries the same compilation mechanism as Python, requiring parsing of all dependencies at once, and does not allow separate compilation. Therefore, different strategies all have a chance to dominate in Java-oriented translations.

### v3

@@ 4. Experimental Design
Category Model Model Type Time Size Training Base (Tokens) In/Out (Tokens)
Commercial DeepSeek-V3 (DeepSeek Team, 2024) Instruction Model 2024-12 671B 14.8 trillion 64k / 4096
Claude-3.5-Sonnet (Cla, [n. d.]) Instruction Model 2024-06 $\sim$175B (Abacha et al., 2024) / 128k / 4096
GPT-4o (Ope, 2025) Instruction Model 2024-05 $\sim$200B (Abacha et al., 2024) / 128k / 4096
General Llama3 (Llama Team, 2024) Instruction Model 2024-04 8B 15.6 trillion 7k / 1024
Llama3 (Llama Team, 2024) Instruction Model 2024-04 70B 15.6 trillion 7k / 1024
Gemma (Team et al., 2024) Base Model 2024-02 7B 6 trillion 11200 Characters/ 1024
Code CodeLlama (Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023) Instruction Model 2023-08 7B 500 billion 8000 Characters/ 1024
CodeGemma (Team, 2024) Base Model 2024-04 7B 500 billion 11200 Characters/ 1024
@@
    In the following sections, “DeepSeek” refers to DeepSeek-V3 and “Claude” indicates Claude-3.5-Sonnet. We also use “Llama3-70B” and “Llama3-8B” to denote the two versions of Llama3.
@@ 4.1. Studied LLMs
Table 2 lists our studied LLMs with their categories, model type, released times (Time), parameter sizes (Size), pretraining token numbers (Training Base), and the input/output context window token limits (In/Out). For those models whose context windows are measured in characters, we specified them separately. As can be seen, to extensively explore the performance of latest LLMs, we select LLMs from general (i.e., Llama 3 (Llama Team, 2024) and Gemma (Team et al., 2024)), code (i.e., CodeLlama (Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023) and CodeGemma (Team, 2024)), and commercial kinds  [^1] We define commercial LLMs as large language models with substantial parameter sizes (typically exceeding 100B), developed by organizations for commercial purposes, and providing publicly accessible interfaces for interactive dialogue and text-based applications. (i.e., DeepSeek-V3 (DeepSeek Team, 2024), GPT-4o (Ope, 2025) and Claude-3.5-sonnet (Cla, [n. d.])), released from late 2023 to late 2024. Furthermore, for LLMs from the same family, we also select two different sizes (i.e., Llama3-8B and Llama3-70B) to study the influence of the volume of model parameters.
@@ 4.4. Implementation Details
We utilize eight LLMs for our experiments. Specifically, we employ the Deepseek, GPT-4o and Claude via their official API interface(Deepseek, 2024; Ope, 2025; Cla, [n. d.]). Additionally, CodeGemma, Gemma, and Llama3-70B are accessed through the open API interface provided by NVIDIA(NVIDIA, 2024), while Llama3-8B and CodeLlama are operated via the open API interface provided by Baidu Qianfan platform(Baidu Qianfan, 2024). The hyperparameters for the code translation process are set as follows: $temperature=0.8$, $top\_p=0.95$, $top\_k=50$, and $n=1$, which means we only fetch LLMs’ first translated candidates for evaluation during inference. All other hyperparameters are kept by default. To mitigate the impact of LLM randomness on result reliability, following prior works(Liu et al., 2024d)(Yan et al., 2023b), we repeat the experiments three times under the same settings and report the averaged results.
Figure 1. Translation results of CSR, CAc and CAm on benchmarks of Yang et al. and ClassEval-T
@@
Gemma 19.15 36.44 89.01 1.06 2.94 9.57 12.41 31.43 78.72 4.26 5.35 18.79 3.90 5.09 26.24 5.32 5.01 9.93
CodeLlama 19.86 39.03 84.40 1.77 3.45 7.45 13.12 32.64 76.95 4.61 5.53 35.11 3.90 6.56 46.81 3.55 5.18 13.12
CodeGemma 25.89 48.70 90.78 6.38 8.12 25.89 18.09 41.62 83.33 10.99 13.30 34.40 7.45 11.14 56.74 8.87 8.89 24.11
Llama3-8B 27.30 50.86 89.01 5.32 6.74 32.98 18.09 46.63 89.72 5.32 7.60 44.33 6.38 10.88 61.70 11.35 11.14 33.69
Llama3-70B 35.46 58.55 90.07 6.38 6.91 32.98 25.89 55.44 94.68 9.22 11.14 71.28 15.6 18.74 77.66 11.35 10.71 30.14
GPT-4o 43.97 64.16 99.65 21.63 24.96 82.62 32.98 60.71 98.94 17.02 17.70 81.21 19.5 21.33 92.91 19.15 24.35 84.75
DeepSeek 45.04 65.98 100.0 24.47 29.79 80.5 35.11 60.97 98.58 26.60 27.37 87.59 22.34 24.18 89.36 20.92 24.53 78.72
Claude 42.20 65.98 99.65 27.30 30.92 85.82 31.56 61.49 98.94 20.21 22.63 87.94 14.54 16.15 91.13 19.86 22.11 86.17
Average 32.36 53.71 92.82 11.79 14.23 44.73 23.41 48.87 89.98 12.28 13.83 57.58 11.70 14.26 67.82 12.55 13.99 45.08
@@
    Table cells with light cyan background denote the highest performance in terms of CAc/CAm/CSR on each translation pair among various LLMs. Bold and underlined values in the Average row represent the highest and second-highest average performance respectively.
@@ 5.1. RQ1: Overall Correctness
Apart from the experiments on ClassEval-T, we also evaluate each studied LLM on Yang et al.’s code translation dataset (Yang et al., 2024a), which is cleaned from (Roziere et al., 2020), one of the most widely used method-level datasets in the code translation field (Roziere et al., 2021; Liu et al., 2023; Szafraniec et al., 2023). Figure 1 demonstrates the correctness comparison between code translation benchmarks of ClassEval-T and Yang et al. among diverse LLMs in terms of average CA and CSR, where CA is added with error bars of 95% confidence interval while CSR is wrapped with ranges among repeated experiments. In particular, the holistic translation strategy is adopted here, and detailed strategy comparison results are showcased in Section 5.2. As can be seen, both class-level CA (CAc) and method-level CA (CAm) are adopted for assessment on ClassEval-T, while Yang et al.’s dataset is solely evaluated on CAm as it only consists of method-level code translation samples. Table 3 presents specific experimental results with the holistic translation strategy on ClassEval-T. Based on Table 3 and Figure 1, we have the following observations.
Method-level code translation v.s. class-level code translation. As shown in Figure 1, the code translation performance of LLMs degenerates significantly and consistently on all translation pairs when converting from method-level to class-level benchmark. Specifically, all LLMs studied can achieve CAm scores over 60% and CSR scores over 80% on almost all translation pairs in Yang et al.’s benchmark. However, when assessing them on the ClassEval-T benchmark, LLMs’ code translation performance declines 51.14%–79.05% on average in terms of CAm while 3.12%–56.24% on average in terms of CSR, where half of the translation results suffer over 80% decline in terms of CAm, and almost half of them suffer over 20% decline in terms of CSR. Besides, based on Yang et al.’s benchmark, it is hard to discriminate which LLM performs better, as method-level code translation benchmarks only contain standalone functions, they are shorter, lack diverse dependencies and complex contextual information, making LLMs easy to handle such samples and demonstrate inapparent performance discrepancies. Even in domain-specific PLs, such as Rust and Swift, a recent study (Tao et al., 2024) also unveiled a similar result. Therefore, we conclude that standalone functions are nearly no longer effective in assessing LLMs’ code translation performance, but existing LLMs still have limited capability in complex coding tasks that bring about brand-new challenges, such as class-level code translation.
@@
Comparison among LLMs. Focusing on the results of the ClassEval-T benchmark, the three commercial LLMs consistently perform the best and are nearly neck-to-neck with each other, due to their much larger number of parameters than other LLMs (Abacha et al., 2024). On average, commercial LLMs achieve a CAc score of 26.91%, a CAm score of 36.96%, and a CSR score of 90.25% across various translation pairs. Regarding the comparison between general and code LLMs, CodeGemma, a code LLM, always performs better than Gemma, a general LLM, on all evaluation metrics among various translation pairs. However, Llama3-8B performs better than CodeLlama in most of the translation pairs. The reason behind this performance difference is that CodeGemma is enhanced from Gemma with an extra 500 billion code tokens for code-specific pretraining, resulting in its superior performance against Gemma. In contrast, CodeLlama is based on Llama2 with an extra 500 billion code samples (1.25 times more than Llama2 (Touvron et al., 2023)) for finetuning, while Llama3-8B was pre-trained on 15 trillion tokens (seven times more than Llama2), with a relatively larger parameter size, leading to its superiority in code translation task against CodeLlama. The above phenomenon and analysis indicate that code LLMs are not necessarily better than general LLMs, even though they are specially pre-trained on code corpus. Delving deep into each category, CodeGemma constantly outperforms CodeLlama, while Llama3-8B constantly outperforms Gemma in terms of both CA and CSR, which are in accordance with evaluation results in other coding tasks conducted by recent studies (Team, 2024; Dubey et al., 2024). In addition, it is obvious that Llama3-70B always generates more samples that can pass test suites than Llama3-8B does among all translation pairs, even though Llama3-8B can generate more compilable samples in certain translation pairs, illustrating that LLMs with more parameters carry more powerful code translation ability, given other factors the same.
@@
Comparison among PLs. Based on Figure 1 and Table 3, LLMs generally perform much better on Python-oriented translations (e.g., C++/Java-to-Python) in terms of all evaluation metrics. However, concerning the rest of the code translation pairs, LLMs’ performance drops dramatically. Specifically, for C++-oriented translations, the CAc decreased by 56.36%, CAm decreased by 72.49%, and CSR decreased by 50.87%. For Java-oriented translations, each evaluation metric decreased by 57%, 72.62%, and 31.40% in order. The performance differences can be attributed to pre-training data biases. Recent LLMs and those included in this paper normally use open-sourced code samples on GitHub for coding enhancement pretraining (Zhao et al., 2024; Liu et al., 2024a; Team, 2024; Roziere, Baptiste and Gehring, Jonas and Gloeckle, Fabian and Sootla, Sten and Gat, Itai and Tan, Xiaoqing Ellen and Adi, Yossi and Liu, Jingyu and Sauvestre, Romain and Remez, Tal and others, 2023; Tao et al., 2024), while Python has achieved the second most used PL on GitHub since 2022, as reported in (GitHub Octoverse, 2024). Therefore, LLMs potentially gained more knowledge of Python generation, which results in its much superior advantages against C++/Java-oriented translations.
@@ 5.2. RQ2: Translation Strategies
Figure 2 demonstrates the experimental results on ClassEval-T with different code translation strategies mentioned in Section 4.2, where each column showcases experimental results of diverse LLMs in terms of each evaluation metric (either CAc, CAm, or CSR) among six translation pairs. All are recorded in their average scores wrapped with ranges among repeated experiments. Our observations are listed below.
Analysis among LLMs. As shown in Figure 2, the holistic translation strategy is always the best selection for commercial LLMs, outperforming the other two strategies by an average of 341.51% in terms of CAc, by 310.62% in terms of CAm, and by 157.64% in terms of CSR, while for other LLMs, it depends on different translation scenarios. One potential reason might be that commercial LLMs carry a large number of parameters (¿100B), which enhances their ability to understand and generate code effectively. Thus, it is not difficult for these models to complete the code translation task given the entire class as inputs in the holistic strategy. However, other LLMs, with smaller parameter sizes, carry relatively limited capability in leveraging long input contexts, as revealed by (Liu et al., 2024b). On the other hand, we observed that, on average, 8.87% of the test samples exceeded the model’s context window length limitation, which in turn constrains the applicability of the holistic translation strategy to some extent. As a result, holistic translation strategy does not show overwhelming advantages against min-dependecy strategy, and in some cases, the min-dependency strategy does indeed yield higher performance. Besides, LLMs with the standalone strategy perform the worst on CAc and CAm in most of the translation scenarios, as we do not provide contextual dependencies but a standalone class skeleton or method during the translation, leading to the translated methods or fields being highly possible to have conflicts with other translated counterparts in classes. For example, one method invoked by another may be translated to different method names during the translation of its caller and itself, causing conflicts in a class. However, this is not the case in terms of CSR. When translating with the standalone strategy, LLMs tend to generate some contextual dependencies based on their imagination; thus, discriminating the translated results from other generated contexts via manually crafted rules is one of our processes for the standalone translation strategy. Considering that the translation process of LLMs is uncontrollable, which means they may rename the source program completely, making our rules difficult to capture the translated results. In this paper, we deem the above a failed translation, and no method will be inserted into the translated class skeleton. This will definitely affect functionality correctness but not the program compilation, leading to an overestimation of CSR among LLMs in the standalone translation strategy.
@@
Analysis among PLs. The selection of translation strategy also relies on target PLs. As shown in Figure 2, the holistic strategy always performs the best in Python-oriented and Java-oriented translations in terms of CAc and CAm, outperforming other strategies on average by 42.05%–84.40% in terms of CAc while by 35.22%–59.09% in terms of CAm. However, focusing on C++-oriented translation, most LLMs with holistic and min-dependency strategies perform neck-to-neck, except for commercial LLMs whose holistic strategy still outperforms but the gap to min-dependecy strategy shrinks. The reasons are multifaceted: (1) Because C++ requires explicit memory allocation and deallocation, lacks rich third-party libraries, and has other language-specific characteristics, implementing the same functionality typically results in longer programs compared to Java and Python. Based on tokenization averaging across all LLMs, C++ code in ClassEval-T has an average of 736.91 tokens, while Python and Java code contains an average of 332.29 and 575.15 tokens. Commercial LLMs have large input and output context windows, so no samples are truncated due to token limitations, allowing the holistic strategy to remain the most optimal. However, for other LLMs, although only an average of 1.43% of samples exceed the LLM’s input limit, a larger challenge arises with C++ data. An expected 19.79% of C++ samples exceeds the output token limit, compared to only 0.64% for Python and 6.17% for Java (To simplify the statistical process, we use the ground truth as the estimated output length). This means that, while most samples can be fully input into the LLM, a significantly higher proportion of C++ samples fail to output completely, reducing the effectiveness of the holistic strategy. As a result, when translating C++, the min-dependency and holistic strategies perform neck-and-neck for LLMs other than commercial LLMs.(2) The min-dependency strategy aligns with C++’s separate compilation feature, enabling flexible code splicing and tolerance for unimplemented sections (Favre, Jean-Marie, 1995). In addition, C++ is more complex than Python and Java, especially in terms of manual memory management, template systems, and strict dependency management. Generating code step by step reduces the information that LLM needs to pay attention to each time, reducing the possibility of LLM generating incorrect code. Therefore, in the case of translating other PLs to C++, the gap between the holistic and min-dependency strategies is smaller for commercial LLMs.